# Equivariant Ehrhart theory, commutative algebra and invariant triangulations of polytopes

### v2

@@ 4.3. Applications to equivariant Ehrhart theory
Let $M=\mathbb{Z}^{3}/\mathbb{Z}(e_{1}+e_{2}+e_{3})\oplus\mathbb{Z}$ and let $\bar{e}_{i}$ denote the image of $e_{i}$ in $M$ for $1\leq i\leq 3$. Let $Q=\operatorname{Conv}\left\{{\bar{e}_{1},\bar{e}_{2},\bar{e}_{3}}\right\}$ and let $P=Q\times[0,e_{4}]$. Let $G=\mathbb{Z}/3\mathbb{Z}$ act on $M$ by cycling the first $3$ coordinates. Then $P$ is a $G$-invariant lattice polytope with $6$ vertices and $h^{*}(P,\rho;t)=1+(1+\chi_{\operatorname{reg}})(t+t^{2})$, where $\chi_{\operatorname{reg}}$ is the regular representation of $G$. Let $\mathcal{S}$ be the pulling refinement of $P$ at $e_{4}$. Then $\mathcal{S}$ has $4$ facets: the simplex $\operatorname{Conv}\left\{{Q,e_{4}}\right\}$ and $G\cdot F$, where $F=\operatorname{Conv}\left\{{[\bar{e}_{1},\bar{e}_{2}]\times[0,e_{4}],e_{4}}\right\}$ is not a simplex. Applying a pulling refinement to $F$ at any vertex of $[\bar{e}_{1},\bar{e}_{2}]\times[0,e_{4}]$ gives a lattice triangulation $\mathcal{T}$ of $F$. Then there is a unique $G$-invariant (non-regular) lattice triangulation of $P$ that contains $\operatorname{Conv}\left\{{Q,e_{4}}\right\}$ and restricts to $g\cdot\mathcal{T}$ on $g\cdot F$ for all $g$ in $G$. On the other hand, by Lemma 4.29, there does not exist $f\in\mathbb{C}[M]$ that is $G$-invariant and nondegenerate with respect to $P$. In particular, by Theorem 1.4, there does not exist a $G$-invariant regular lattice triangulation of $P$. Let $x_{1},x_{2},x_{3},x_{4}$ be the degree $1$ variables in $S_{P}$ corresponding to the lattice points $\bar{e}_{1},\bar{e}_{2},\bar{e}_{3},0$ of $Q$. Let $y_{1},y_{2},y_{3},y_{4}$ denote the degree $1$ variables in $S_{P}$ corresponding to the lattice points $\bar{e}_{1}+e_{4},\bar{e}_{2}+e_{4},\bar{e}_{3}+e_{4},e_{4}$ of $Q\times\{e_{4}\}$. If $\sigma$ is a generator of $G$, let $\theta_{1}=x_{1}+x_{2}+2x_{3}+y_{1}+2y_{2}+y_{3}$, $\theta_{2}=\sigma\cdot\theta_{1}$, $\theta_{3}=\sigma^{2}\cdot\theta_{1}$ and $\theta_{4}=x_{4}-y_{4}$. Then one may verify using Macaulay2 that $\theta_{1},\theta_{2},\theta_{3},\theta_{4}$ is a l.s.o.p. for $S_{P}$. Also, $[\mathbb{C}\theta_{1}+\cdots+\mathbb{C}\theta_{4}]=[\widetilde{M}_{\mathbb{C}}]=1+\chi_{\operatorname{reg}}$ in $R(G)$.
@@
We continue with Example 3.22. Recall that $G=\mathbb{Z}/5\mathbb{Z}=\langle\sigma\rangle$ and $P$ is a $4$-dimensional reflexive polytope with 10 vertices in two orbits $\{\bar{e}_{i}:1\leq i\leq 5\}$ and $\{\bar{f}_{i}:1\leq i\leq 5\}$, and that $P$ admits a non-regular $G$-invariant lattice triangulation. Recall that $P$ has a face $Q=\operatorname{Conv}\left\{{\bar{e}_{1},\bar{e}_{3},\bar{f}_{1},\bar{f}_{2}}\right\}\cong[0,1]^{2}$. Then Lemma 4.29 implies that there does not exist $f\in\mathbb{C}[M]$ that is $G$-invariant and nondegenerate with respect to $P$. Let $x_{i}$ and $y_{i}$ denote the degree $1$ variables in $S_{P}$ corresponding to the lattice points $\bar{e}_{i}$ and $\bar{f}_{i}$ respectively for $1\leq i\leq 5$. Let $\theta_{i}=\sigma^{i-1}\cdot(x_{1}+y_{2})$ for $1\leq i\leq 5$. Then one may verify using Macaulay2 that $\theta_{1},\ldots,\theta_{5}$ is a l.s.o.p. for $S_{P}$. Also, $[\mathbb{C}\theta_{1}+\cdots+\mathbb{C}\theta_{5}]=[\widetilde{M}_{\mathbb{C}}]=\chi_{\operatorname{reg}}$ in $R(G)$, where $\chi_{\operatorname{reg}}$ is the regular representation of $G$.
@@
Recall that a signed permutation in $n$ variables is a permutation $\pi$ of {±$\{\pm 1,\ldots,\pm n\}$ such that $\pi(-i)=-\pi(i)$ for $1\leq i\leq n$. Let $B_{3}$ be the group of signed permuations in $3$ variables. Let $G=\{\operatorname{id},\sigma,\tau,\sigma\tau\}\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ be the subgroup of $B_{3}$ with elements $\sigma=(12)(-3)$, $\tau=(1-2)(-3)$ and $\sigma\tau=\tau\sigma=(-1)(-2)$. Here we use cycle notation so, for example, $\tau$ has orbits $\{1,-2\}$,$\{-1,2\}$ and $\{3,-3\}$. Consider the affine lattice $M^{\prime}=\{(u_{1},u_{2},u_{3})\in\mathbb{Z}^{3}:u_{1}+u_{2}+u_{3}=1\mod 2\}$ in $\mathbb{R}^{3}$ with the corresponding action of $B_{3}$, and let $P=[-1,1]^{3}$. Then $P$ is $B_{3}$-invariant and is a ‘unit cube’ with respect to the lattice $M^{\prime}$. Geometrically, $\sigma$, $\tau$ and $\sigma\tau$ correspond to $180$ degree rotations through the lines $\{x=y,z=0\}$, $\{x=-y,z=0\}$ and $\{x=y=0\}$ respectively. See [EKS23, Figure 12] for a beautiful illustration. There are two $G$-orbits of lattice points in $P$: $\{(1,1,1),(1,1,-1),(-1,-1,1),(-1,-1,-1)\}$ and $\{(1,-1,1),(1,-1,-1),(-1,1,1),(-1,1,-1)\}$. Also, $h^{*}(P,\rho;t)=1+\chi_{\operatorname{reg}}t+t^{2}$, where $\chi_{\operatorname{reg}}$ is the regular representation of $G$ (see the proof of [EKS23, Theorem 1.2]). Then [EKS23, Theorem 1.2] implies that there does not exist a $G$-invariant polynomial that is nondegenerate with respect to $P$. This also follows from Lemma 4.29 applied to the face $Q$ of $P$ with vertices $(1,1,1),(1,1,-1),(1,-1,1),(1,-1,-1)$. In particular, by Theorem 1.4, $P$ does not admit a $G$-invariant regular lattice triangulation. Since all lattice triangulations of the unit cube are regular [DLRS10, Theorem 6.3.10], we deduce that $P$ does not admit a $G$-invariant lattice triangulation. On the other hand, let $x_{i,j,k}$ denote the degree $1$ variable in $S_{P}$ corresponding to the vertex $(i,j,k)$ of $P$. Fix a nonzero $\lambda\in\mathbb{C}$ with $\lambda^{4}\neq 1$. Let
@@
If we let $a:=x_{-1,-1,1},b:=x_{-1,-1,-1},c:=x_{1,1,-1},d:=x_{1,1,1}$, then one may compute that $S_{P}/(\theta_{1},\theta_{2},\theta_{3},\theta_{4})$ is isomorphic to the polynomial ring in the variables $a,b,c,d$, subject to the relations $a^{2}=b^{2}=c^{2}=d^{2}=ad=cb=0$ and $ab=cd=\lambda^{2}ac=\lambda^{2}bd$. In particular, $\theta_{1},\theta_{2},\theta_{3},\theta_{4}$ is a l.s.o.p. for $S_{P}$. Also, $[\mathbb{C}\theta_{1}+\cdots+\mathbb{C}\theta_{4}]=[\widetilde{M}_{\mathbb{C}}^{\prime}]=\chi_{\operatorname{reg}}$ in $R(G)$.
@@
Let $\mathcal{S}$ be a $G$-invariant lattice triangulation of $P$. Recall that the fan over the faces of $\mathcal{S}$ is the fan $\Sigma_{\mathcal{S}}=\{C_{F}:F\in\mathcal{S}\}$ that refines $C_{P}$. Let $\operatorname{Box}(\mathcal{S}):=\bigcup_{F\in\mathcal{S}}\operatorname{Box}^{\circ}(C_{F})$. Then $G$ acts permuting the elements of $\operatorname{Box}(\mathcal{S})$. Let $\operatorname{Box}(\mathcal{S})/G$ denote the set of $G$-orbits of $\operatorname{Box}(\mathcal{S})$. For $u\in\operatorname{Box}^{\circ}(C_{F})$, write $C_{u}:=C_{F}$ i.e. $C_{u}$ is the smallest cone in $\Sigma_{\mathcal{S}}$ containing $u$, and let $G_{u}$ be the stablizer of $C_{u}$ in $G$. Also, let $[u]\in\operatorname{Box}(\mathcal{S})/G$ denote the $G$-orbit of $u$. Recall that $\operatorname{pr}:\widetilde{M}\to\mathbb{Z}$ denotes projection onto the last coordinate. Given a subgroup $H$ of $G$ and an element $r\in R(H)$, there is well-defined element $\operatorname{Ind}_{H}^{G}r\in R(G)$ such that if $r=[W]$ for a $\mathbb{C}H$-module $W$, then $\operatorname{Ind}_{H}^{G}r=[\operatorname{Ind}_{H}^{G}W]$. If $r(t)=\sum_{m\geq 0}r_{m}t^{m}\in R(H)[[t]]$ then we write $\operatorname{Ind}_{H}^{G}r(t):=\sum_{m\geq 0}(\operatorname{Ind}_{H}^{G}r_{m})t^{m}\in R(G)[[t]]$. We now prove the following formula for $h^{*}(P,\rho;t)$ that generalizes known formulas when $P$ is a simplex [Sta11a, Proposition 6.1] or when $\Sigma_{\mathcal{S}}$ is a smooth fan i.e. $\mathcal{S}$ is unimodular [Sta11a, Proposition 8.1]. The case when $G$ is trivial is due to Betke and McMullen [BM85, Theorem 1]. Geometrically, the proposition below states that $h^{*}(P,\rho;t)$ is the equivariant Hilbert polynomial of the orbifold cohomology ring of the toric variety $X_{\Sigma_{\mathcal{S}}}$ associated to $\Sigma_{\mathcal{S}}$ (see, for example, [Kar08, (1.1)]).
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensonal lattice polytope. Suppose there exists a $G$-invariant lattice triangulation $\mathcal{S}$ of $P$. Then, with the notation above,
@@
On the right hand side, for any $u\in\operatorname{Box}^{\circ}(C_{F})$, $\bigoplus_{u^{\prime}\in[u]}x^{u^{\prime}}\overline{\operatorname{SR}}(\operatorname{lk}_{\Sigma_{\mathcal{S}}}(C_{u^{\prime}}))$ is isomorphic as a graded $\mathbb{C}G$-module to the induced representation $\operatorname{Ind}_{G_{u}}^{G}x^{u}\overline{\operatorname{SR}}(\operatorname{lk}_{\Sigma_{\mathcal{S}}}(C_{u}))$. The result follows by taking equivariant Hilbert series of both sides of (15). ∎

### v3

@@ 4.3. Applications to equivariant Ehrhart theory
Let $M=\mathbb{Z}^{3}/\mathbb{Z}(e_{1}+e_{2}+e_{3})\oplus\mathbb{Z}$ and let $\bar{e}_{i}$ denote the image of $e_{i}$ in $M$ for $1\leq i\leq 3$. Let $Q=\operatorname{Conv}\left\{{\bar{e}_{1},\bar{e}_{2},\bar{e}_{3}}\right\}$ and let $P=Q\times[0,e_{4}]$. Let $G=\mathbb{Z}/3\mathbb{Z}$ act on $M$ by cycling the first $3$ coordinates. Then $P$ is a $G$-invariant lattice polytope with $6$ vertices and $h^{*}(P,\rho;t)=1+(1+\chi_{\operatorname{reg}})(t+t^{2})$, where $\chi_{\operatorname{reg}}$ is the regular representation of $G$. Let $\mathcal{S}$ be the pulling refinement of $P$ at $e_{4}$. Then $\mathcal{S}$ has $4$ facets: the simplex $\operatorname{Conv}\left\{{Q,e_{4}}\right\}$ and $G\cdot F$, where $F=\operatorname{Conv}\left\{{[\bar{e}_{1},\bar{e}_{2}]\times[0,e_{4}],e_{4}}\right\}$ is not a simplex. Applying a pulling refinement to $F$ at any vertex of $\bar{e}_{1},\bar{e}_{2}]\times[0,e_{4}]$ gives a lattice triangulation $\mathcal{T}$ of $F$. Then there is a unique $G$-invariant (non-regular) lattice triangulation of $P$ that contains $\operatorname{Conv}\left\{{Q,e_{4}}\right\}$ and restricts to $g\cdot\mathcal{T}$ on $g\cdot F$ for all $g$ in $G$. On the other hand, by Lemma 4.29, there does not exist $f\in\mathbb{C}[M]$ that is $G$-invariant and nondegenerate with respect to $P$. In particular, by Theorem 1.4, there does not exist a $G$-invariant regular lattice triangulation of $P$.
Let $x_{1},x_{2},x_{3},x_{4}$ be the degree $1$ variables in $S_{P}$ corresponding to the lattice points $\bar{e}_{1},\bar{e}_{2},\bar{e}_{3},0$ of $Q$. Let $y_{1},y_{2},y_{3},y_{4}$ denote the degree $1$ variables in $S_{P}$ corresponding to the lattice points $\bar{e}_{1}+e_{4},\bar{e}_{2}+e_{4},\bar{e}_{3}+e_{4},e_{4}$ of $Q\times\{e_{4}\}$. If $\sigma$ is a generator of $G$, let $\theta_{1}=x_{1}+x_{2}+2x_{3}+y_{1}+2y_{2}+y_{3}$, $\theta_{2}=\sigma\cdot\theta_{1}$, $\theta_{3}=\sigma^{2}\cdot\theta_{1}$ and $\theta_{4}=x_{4}-y_{4}$. Then one may verify using Macaulay2 that $\theta_{1},\theta_{2},\theta_{3},\theta_{4}$ is a l.s.o.p. for $S_{P}$. Also, $\mathbb{C}\theta_{1}+\cdots+\mathbb{C}\theta_{4}]=[\widetilde{M}_{\mathbb{C}}]=1+\chi_{\operatorname{reg}}$ in $R(G)$.
@@
We continue with Example 3.23. Recall that $G=\mathbb{Z}/5\mathbb{Z}=\langle\sigma\rangle$ and $P$ is a $4$-dimensional reflexive polytope with 10 vertices in two orbits $\{\bar{e}_{i}:1\leq i\leq 5\}$ and $\{\bar{f}_{i}:1\leq i\leq 5\}$, and that $P$ admits a non-regular $G$-invariant lattice triangulation. Recall that $P$ has a face $Q=\operatorname{Conv}\left\{{\bar{e}_{1},\bar{e}_{3},\bar{f}_{1},\bar{f}_{2}}\right\}\cong[0,1]^{2}$. Then Lemma 4.29 implies that there does not exist $f\in\mathbb{C}[M]$ that is $G$-invariant and nondegenerate with respect to $P$. Let $x_{i}$ and $y_{i}$ denote the degree $1$ variables in $S_{P}$ corresponding to the lattice points $\bar{e}_{i}$ and $\bar{f}_{i}$ respectively for $1\leq i\leq 5$. Let $\theta_{i}=\sigma^{i-1}\cdot(x_{1}+y_{2})$ for $1\leq i\leq 5$. Then one may verify using Macaulay2 that $\theta_{1},\ldots,\theta_{5}$ is a l.s.o.p. for $S_{P}$. Also, $\mathbb{C}\theta_{1}+\cdots+\mathbb{C}\theta_{5}]=[\widetilde{M}_{\mathbb{C}}]=\chi_{\operatorname{reg}}$ in $R(G)$, where $\chi_{\operatorname{reg}}$ is the regular representation of $G$.
@@
Recall that a signed permutation in $n$ variables is a permutation $\pi$ of $\{\pm 1,\ldots,\pm n\}$ such that $\pi(-i)=-\pi(i)$ for $1\leq i\leq n$. Let $B_{3}$ be the group of signed permuations in $3$ variables. Let $G=\{\operatorname{id},\sigma,\tau,\sigma\tau\}\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ be the subgroup of $B_{3}$ with elements $\sigma=(12)(-3)$, $\tau=(1-2)(-3)$ and $\sigma\tau=\tau\sigma=(-1)(-2)$. Here we use cycle notation so, for example, $\tau$ has orbits $\{1,-2\}$$\{-1,2\}$ and $\{3,-3\}$. Consider the affine lattice $M^{\prime}=\{(u_{1},u_{2},u_{3})\in\mathbb{Z}^{3}:u_{1}+u_{2}+u_{3}=1\mod 2\}$ in $\mathbb{R}^{3}$ with the corresponding action of $B_{3}$, and let $P=[-1,1]^{3}$. Then $P$ is $B_{3}$-invariant and is a ‘unit cube’ with respect to the lattice $M^{\prime}$. Geometrically, $\sigma$, $\tau$ and $\sigma\tau$ correspond to $180$ degree rotations through the lines $\{x=y,z=0\}$, $\{x=-y,z=0\}$ and $\{x=y=0\}$ respectively. See [EKS23, Figure 12] for a beautiful illustration. There are two $G$-orbits of lattice points in $P$: $\{(1,1,1),(1,1,-1),(-1,-1,1),(-1,-1,-1)\}$ and $\{(1,-1,1),(1,-1,-1),(-1,1,1),(-1,1,-1)\}$. Also, $h^{*}(P,\rho;t)=1+\chi_{\operatorname{reg}}t+t^{2}$, where $\chi_{\operatorname{reg}}$ is the regular representation of $G$ (see the proof of [EKS23, Theorem 1.2]).
Then [EKS23, Theorem 1.2] implies that there does not exist a $G$-invariant polynomial that is nondegenerate with respect to $P$. This also follows from Lemma 4.29 applied to the face $Q$ of $P$ with vertices (1,1,1),(1,1,1),(1,1,1),(1,1,1)(1,1,1),(1,1,-1),(1,-1,1),(1,-1,-1). In particular, by Theorem 1.4, $P$ does not admit a $G$-invariant regular lattice triangulation. Since all lattice triangulations of the unit cube are regular [DLRS10, Theorem 6.3.10], we deduce that $P$ does not admit a $G$-invariant lattice triangulation.
On the other hand, let $x_{i,j,k}$ denote the degree $1$ variable in $S_{P}$ corresponding to the vertex $(i,j,k)$ of $P$. Fix a nonzero $\lambda\in\mathbb{C}$ with $\lambda^{4}\neq 1$. Let
@@
If we let $a:=x_{-1,-1,1},b:=x_{-1,-1,-1},c:=x_{1,1,-1},d:=x_{1,1,1}$, then one may compute that $S_{P}/(\theta_{1},\theta_{2},\theta_{3},\theta_{4})$ is isomorphic to the polynomial ring in the variables $a,b,c,d$, subject to the relations $a^{2}=b^{2}=c^{2}=d^{2}=ad=cb=0$ and $ab=cd=\lambda^{2}ac=\lambda^{2}bd$. In particular, $\theta_{1},\theta_{2},\theta_{3},\theta_{4}$ is a l.s.o.p. for $S_{P}$. Also, $\mathbb{C}\theta_{1}+\cdots+\mathbb{C}\theta_{4}]=[\widetilde{M}_{\mathbb{C}}^{\prime}]=\chi_{\operatorname{reg}}$ in $R(G)$.
@@
Let $\mathcal{S}$ be a $G$-invariant lattice triangulation of $P$. Recall that the fan over the faces of $\mathcal{S}$ is the fan $\Sigma_{\mathcal{S}}=\{C_{F}:F\in\mathcal{S}\}$ that refines $C_{P}$. Let $\operatorname{Box}(\mathcal{S}):=\bigcup_{F\in\mathcal{S}}\operatorname{Box}^{\circ}(C_{F})$. Then $G$ acts permuting the elements of $\operatorname{Box}(\mathcal{S})$. Let $\operatorname{Box}(\mathcal{S})/G$ denote the set of $G$-orbits of $\operatorname{Box}(\mathcal{S})$. For $u\in\operatorname{Box}^{\circ}(C_{F})$, write $C_{u}:=C_{F}$ i.e. $C_{u}$ is the smallest cone in $\Sigma_{\mathcal{S}}$ containing $u$, and let $G_{u}$ be the stablizer of $C_{u}$ in $G$. Also, let $u]\in\operatorname{Box}(\mathcal{S})/G$ denote the $G$-orbit of $u$. Recall that $\operatorname{pr}:\widetilde{M}\to\mathbb{Z}$ denotes projection onto the last coordinate. Given a subgroup $H$ of $G$ and an element $r\in R(H)$, there is well-defined element $\operatorname{Ind}_{H}^{G}r\in R(G)$ such that if $r=[W]$ for a $\mathbb{C}H$-module $W$, then $\operatorname{Ind}_{H}^{G}r=[\operatorname{Ind}_{H}^{G}W]$. If $r(t)=\sum_{m\geq 0}r_{m}t^{m}\in R(H)[[t]]$ then we write $\operatorname{Ind}_{H}^{G}r(t):=\sum_{m\geq 0}(\operatorname{Ind}_{H}^{G}r_{m})t^{m}\in R(G)[[t]]$. We now prove the following formula for $h^{*}(P,\rho;t)$ that generalizes known formulas when $P$ is a simplex [Sta11a, Proposition 6.1] or when $\Sigma_{\mathcal{S}}$ is a smooth fan i.e. $\mathcal{S}$ is unimodular [Sta11a, Proposition 8.1]. The case when $G$ is trivial is due to Betke and McMullen [BM85, Theorem 1]. Geometrically, the proposition below states that $h^{*}(P,\rho;t)$ is the equivariant Hilbert polynomial of the orbifold cohomology ring of the toric variety $X_{\Sigma_{\mathcal{S}}}$ associated to $\Sigma_{\mathcal{S}}$ (see, for example, [Kar08, (1.1)]).
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensional lattice polytope. Suppose there exists a $G$-invariant lattice triangulation $\mathcal{S}$ of $P$. Then, with the notation above,
@@
[…]